OpenStudy (anonymous):

Area of a rectangular playground is 20 squared yards. The length is 5 times longer, than its width. Find the length and width of the playground.

Parth (parthkohli):

Well so basically length is 5x and width is x. area = l * w, so \(\Large \color{Black}{\Rightarrow 5x \times x = 20 }\) \(\Large \color{Black}{\Rightarrow 5x^2 = 20 }\) \(\Large \color{Black}{\Rightarrow x^2 = 4 }\) \(\Large \color{Black}{\Rightarrow x = 2 }\) So the width is x = 2 length is 2 * 5 = 10

OpenStudy (anonymous):

can you also help me solve an eqution?

OpenStudy (anonymous):

equation*

Parth (parthkohli):

Sure. Yes.

OpenStudy (anonymous):

x^2 + 3 = 9x i know it's going to be; x^2 - 9x + 3 = 0

OpenStudy (anonymous):

then; a= 1 b= -9 c= 3

Parth (parthkohli):

Hmm correct

Parth (parthkohli):

Use the quadratic formula.

Parth (parthkohli):

Do you know what the formula is?

Parth (parthkohli):

It is: \(\Large \color{Black}{\Rightarrow x = {-b \pm \sqrt{b^2 - 4ac} \over 2a} }\)

OpenStudy (anonymous):

\[x = -b \pm \sqrt{b ^{2} - 4ac} / 2a\]

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

i got 13.15 & 4.85

OpenStudy (anonymous):

but it doesn't equal 0

Parth (parthkohli):

Wolfram in this case :)

OpenStudy (anonymous):

lol what? :)

OpenStudy (anonymous):

so what is the answer? im confused lol